# When ControlNet Meets Inexplicit Masks: A Case Study of ControlNet on its Contour-following Ability

### v2

@@ Appendix B Future Works
By far, we have demonstrated the effectiveness of our method in improving ControlNet in dealing with inexplicit masks, which is an essential topic in applying ControlNet in practical usage. Moreover, we have noted that the dramatic performance degradation caused by inexplicit masks is a common issue among methods that inject spatially aligned control signals for spatially controllable T2I generation. The results of T2I-Adapter Mou et al. [2023] are shown in Fig. 13, where large deterioration in the control masks would pose challenges in correctly understanding the spatial control signals. Since our method takes no assumption about the network that injects control signals into the generation process, it has the potential to extend to other adapter-based methods like T2I-Adapter to avoid deviation of spatial control brought by inexplicit masks. We leave this as our future work.

### v3

@@ Appendix E Future Works
By far, we have demonstrated the effectiveness of our method in improving ControlNet in dealing with inexplicit conditions, which is an essential topic in applying ControlNet to practical usage. Moreover, we have noted that the dramatic performance degradation caused by inexplicit conditions is a common issue among methods that inject spatially aligned control signals for spatially controllable T2I generation. The results of T2I-Adapter (Mou et al., 2024) are shown in Fig. 17, where large deterioration in the control masks would pose challenges in correctly understanding the spatial control signals. Since our method takes no assumption about the network that injects control signals into the generation process, it has the potential to extend to other adapter-based methods like T2I-Adapter to avoid deviation of spatial control brought by inexplicit conditions of varying deterioration degrees. We leave this as our future work.
Figure 17. More examples of ControlNet and T2I-adapter illustrating spatially controllable generation with deteriorated masks of varying degrees. Both methods exhibit strong preservation of contours during the generation process. Green denotes the recalled contours and blue denotes the missing ones. Moreover, inexplicit masks with inaccurate contours would cause severe degradation of image fidelity and realism.
Figure 18. Visualized examples to illustrate the inductive bias, where ControlNet-$\boldsymbol{m}_{r}$ implicitly learns the dilated radius $r$ from the data. Therefore, it always assumes a dilation in the provided mask. The ControlNet-$\boldsymbol{m}_{r}$ conditioned on $\boldsymbol{m}_{r}$ (in red boxes) produces objects with similar shapes and sizes of mask $\boldsymbol{m}_{0}$. While ControlNet-$\boldsymbol{m}_{\infty}$ also breaks the inductive bias, our method surpasses it in terms of image fidelity and additional control over the shape priors.
Figure 19. Visualized examples under different hyperparameter settings, i.e., CFG scale, conditioning scale, and condition injection strategy. Though the CFG scale shows little impact on the contour-following ability, reducing conditioning scales and discarding conditions at early reverse sampling stages help to relieve contour instructions, resulting in better results with inexplicit conditional masks. However, it is still tricky and hard to achieve satisfactory results through hyperparameter tuning, especially for deteriorated control masks.
Figure 20. More examples of our Shape-aware ControlNet compared with the vanilla ControlNet, i.e., ControlNet-$\boldsymbol{m}_{0}$ and ControlNet-$\boldsymbol{m}_{\infty}$ given the conditional mask $\boldsymbol{m}_{r}$. Our method exhibits robust performance on inexplicit masks of varying deterioration degrees.
Figure 21. Error analysis of our deterioration estimator at different dilation radius $r$. The overall average $L1$ error is 5.47%. Notably, such errors show little impact on the performance of our Shape-aware ControlNet, referring to our main paper § 6.3.
Figure 22. More examples of our method on TikZ sketches compared with the vanilla ControlNet. Though our Shape-aware ControlNet is only trained with dilated masks, it generalizes well to abstract programmatic masks with inaccurate contours and exhibits advanced performance.
Figure 23. More examples of our method on human scribbles compared with the vanilla ControlNet. Our Shape-aware ControlNet exhibits advanced and robust performance on realistic user-provided masks with inaccurate contours.
Figure 24. More examples of controlling object shape via the deterioration ratio $\rho$. The value of $\rho$ is depicted on the top row. Despite training with $\rho\in[0,1]$, we empirically find it generalizes well to other values and offers additional control over the shape of generated objects. A small $|\rho|$ encourages generated objects to adhere more tightly to the outlines, and vice versa. Moreover, note that adjusting $\rho$ has little impact on image fidelity.
Figure 25. More examples of composable shape-controllable generation with our method. Leveraging a Multi-ControlNet structure (Zhang et al., 2023a), we can assign different priors to each part of the control masks. This enables strict shape control over specific masks, while simultaneously allowing T2I diffusion models to unleash creativity in imagining objects of diverse shapes within inexplicit masks.